# Higher-dimensional flying wing Steady Ricci Solitons

### v1

@@ 5.2. Preservation of symmetry group
In particular, the maximum principle shows that $|X_{i}|_{g(t)}\leq\sup_{\mathbb{S}^{n}}|X_{0}|_{g(t_{i})}$ on $t_{i},T]$. Since the Ricci flow is non-increasing using $\text{\rm Ric}\geq 0$, it follows from the construction that $g(t_{i})\leq Cg_{0}$ almost everywhere. Since $g_{0}$ is a $L^{\infty}$ metric on $M^{n}$, the conclusion follows. ∎
@@
We apply a trick in [8, Lemma 3.2] that if $\varphi$ denotes either the negative part of the lowest eigenvalues $\eta_{-}$ or the positive part of the highest eigenvalues $\eta_{+}$, then it satisfies $\left(\frac{\partial}{\partial t}-\Delta_{g(t)}\right)\varphi\leq R\cdot\varphi$ in the sense of barrier. Particularly, we have
@@
By Theorem 5.2, the Ricci flow smoothing of $g_{0,{\beta}}$ exists. By [19, Theorem 1.6], we might assume $g(t)$ converges to $g_{0,{\beta}}$ as $t\to 0$, outside a set of upper Minkowski dimension $\leq n-2$. By Proposition 5.4, the Ricci flow smoothing preserved the symmetry from the initial singular metrics. By choosing $t$ sufficiently small, the conclusion follows from Hamilton-Perelman distance distortion, for instances see [43, Lemma 3.1]. The family $g_{\beta}(t)$ is continuous by (iii) and Lemma 4.7, using the fact that the push-forward of Ricci flow by the Ricci-harmonic map heat flow is a solution to the Ricci-DeTurck flow for $n-1\geq 3$. For n1=2n-1=2, the higher order stability follows from the explicit ODE construction in [28]. ∎
@@ 6. Steady solitons with prescribed eigenvalues
In this section, we prove our main Theorem. This immediately implies Theorem 1.2 and Theorem 1.3.
@@
So for any $\mathbf{y}\in\Delta_{k}$, for all sufficiently large $j$ we can find $\mathbf{x}_{j}\in\Omega_{k}$ such that $\mathbf{y}=\lim_{j\to\infty}\Lambda_{j}^{*}(\mathbf{x}_{j})$. By the same argument as in [28, Lemma 2.3], after passing to a subsequence, the expanding solitons $\mathcal{E}_{j}(\mathbf{x}_{j})$ converge to a steady gradient soliton with $\text{\rm Rm}\geq 0$ and $R=1$ at the critical point as $j\to\infty$. Moreover, the eigenvalues of the steady soliton at the critical point are given by the vector $\mathbf{y}$. This proves the main theorem. ∎
@@
When $n=5$, by the holonomy classification of Berger and a result of Deng-Zhu [16], the 2-parameter family of $O(3)$-symmetric non-collapsed steadies in Theorem 1.2 have positive curvature operator.
@@ Appendix B Warped product metric and curvature
Let $\{e_{i}\}_{i=0}^{m}$ be a local orthonormal frame on $(0,\pi)\times N$ with respect to $g$ such that $e_{0}=\partial_{r}$. Denote its dual frame by $\{e_{i}^{*}\}_{i=0}^{m}$. By the curvature formula of warped product metric [36, Appendix A], for $i,j,k,l\geq 1$,

### v2

@@ 5.2. Preservation of symmetry group
In particular, the maximum principle shows that $|X_{i}|_{g(t)}\leq\sup_{M}|X_{0}|_{g(t_{i})}$ on $t_{i},T]$. Since the Ricci flow is non-increasing using $\text{\rm Ric}\geq 0$, it follows from the construction that $g(t_{i})\leq Cg_{0}$ almost everywhere. Since $g_{0}$ is a $L^{\infty}$ metric on $M^{n}$, the conclusion follows. ∎
@@
We apply a trick in [10, Lemma 3.2] that if $\varphi$ denotes either the negative part of the lowest eigenvalues $\eta_{-}$ or the positive part of the highest eigenvalues $\eta_{+}$, then it satisfies $\left(\frac{\partial}{\partial t}-\Delta_{g(t)}\right)\varphi\leq R\cdot\varphi$ in the sense of barrier. Particularly, we have
@@
By Theorem 5.2, the Ricci flow smoothing of $g_{0,{\beta}}$ exists. By [21, Theorem 1.6], we might assume $g(t)$ converges to $g_{0,{\beta}}$ as $t\to 0$, outside a set of upper Minkowski dimension $\leq n-2$. By Proposition 5.5, the Ricci flow smoothing preserved the symmetry from the initial singular metrics. By choosing $t$ sufficiently small, the conclusion follows from Hamilton-Perelman distance distortion, for instances see [45, Lemma 3.1]. The family $g_{\beta}(t)$ is continuous by (iii) and Lemma 4.7, using the fact that the push-forward of Ricci flow by the Ricci-harmonic map heat flow is a solution to the Ricci-DeTurck flow for $n-1\geq 3$. For n1=2n-1=2, the higher order stability follows from the explicit ODE construction in [30]. ∎
@@ 6. Steady solitons with prescribed eigenvalues
In this section, we prove our main Theorem. This immediately implies Theorem 1.2.
@@
So for any $\mathbf{y}\in\Delta_{k}$, for all sufficiently large $j$ we can find $\mathbf{x}_{j}\in\Omega_{k}$ such that $\mathbf{y}=\lim_{j\to\infty}\Lambda_{j}^{*}(\mathbf{x}_{j})$. By the same argument as in [30, Lemma 2.3], after passing to a subsequence, the expanding solitons $\mathcal{E}_{j}(\mathbf{x}_{j})$ converge to a steady gradient soliton with $\text{\rm Rm}\geq 0$ and $R=1$ at the critical point as $j\to\infty$. Moreover, the eigenvalues of the steady soliton at the critical point are given by the vector $\mathbf{y}$. This proves the main theorem. ∎
@@
When $n=5$, by Berger’s holonomy classification and a result of Deng-Zhu [18], one may choose $\lambda_{1}>0$ (i.e. $\text{\rm Ric}>0$) so that the 2-parameter family of $O(3)$-symmetric non-collapsed steady solitons in Theorem 1.2 has positive curvature operator. More generally, Theorem 1.1 provides a wealth of new examples of positively curved steady solitons in higher dimensions. For instances, we can consider the examples from Theorem 1.1 with $0<\lambda_{1}<\lambda_{2}<\dots<\lambda_{k}<\lambda_{k+1}=\cdots=\lambda_{n}$ and the following symmetry
$\underbrace{O(1)\times\cdots\times O(1)}_{\text{$k$ times}}\times O(n-k),$
where $n\geq k+2$. With such a choice of $\lambda_{i}$, it follows from the symmetry that there exists a new $2$-parameter family of $O(1)\times O(1)\times O(n-2)$ symmetric steady solitons with $\text{\rm Rm}>0$ and 3 distinct eigenvalues of Ric at the critical point of $f$, for each $n\geq 4$. Similarly by [32], Theorem 1.1 also gives a new $3$-parameter family of $O(1)\times O(1)\times O(1)\times O(n-3)$ symmetric steady solitons with $\text{\rm Rm}>0$ and $4$ distinct eigenvalues of Ric at the critical point of $f$, for each $n\geq 5$. Moreover, the $O(1)\times O(1)\times O(n-2)$ and $O(1)\times O(1)\times O(1)\times O(n-3)$ symmetric examples are noncollapsed when $n\geq 5$ and $n\geq 6$ respectively.
@@ Appendix B Warped product metric and curvature
Let $\{e_{i}\}_{i=0}^{m}$ be a local orthonormal frame on $(0,\pi)\times N$ with respect to $g$ such that $e_{0}=\partial_{r}$. Denote its dual frame by $\{e_{i}^{*}\}_{i=0}^{m}$. By the curvature formula of warped product metric [38, Appendix A], for $i,j,k,l\geq 1$,